Operations Report – May 15th

Crew 212 Operations Report 15-MAY-2019

SOL: 10

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None.

Notes on non-nominal systems: Nothing to report.

Generator (hours run last night): 12:00

On (last night): 7:00 PM

Off (this morning): 7:00 AM

Solar— SOC 74% at 6:15 PM

Diesel Reading – 85%

Propane Reading – 45%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 260 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 122 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143678.1

Toilet tank emptied: No

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes [USED BY DIRECTOR FOR MAINTENANCE]

Hours:

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Opportunity rover used: Yes [USED BY DIRECTOR  FOR MAINTENANCE]

Hours:

Beginning charge: 99%

Ending charge: 74%

Currently charging: No

Curiosity rover used: No

Hours: 96.6h

Beginning charge:

Ending charge:

Currently charging: Yes

Notes on rovers: [From Director] cleaned all the batteries on Spirit today and checked for loose connections. The connections were very corroded but not loose. Then I checked all the cells on the batteries. They were basically fine and I added 1 gallon of distilled water to them. However, there were three cells on one battery that could not be opened unless a security bar was removed. Apparently when others had filled the batteries they thought it would be fine to just ignore those cells rather than bother with the bar. The cells were completely dry. I filled them with about 1/2 gallon of distilled water and secured the bar again. I am currently charging the rover and we will see if that battery can recover. If not, I will switch it out with one on Opportunity for the remainder of the field season.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? Off Campus

General notes and comments:

· The tests run this morning on EVA#11 proved unsuccessful.

· The day was spent working on experiments and doing some fixes in the Station.

Summary of internet: Nothing to report.

Summary of suits and radios: A headset was damaged during EVA#10.

Summary of Hab operations:

· We assembled the new pump in the loft and changed the water filter. The area was dried and cleaned, and the new pump has been fixed to the plywood. The new pump maintains the pressure higher than before, and this fixed the heater problem; now we have enough pressure to activate the hot water. The leak in the commander’s room is stopped, and humidity is reduced, but the damage done by the old pump will require a remaking of the ceiling.

· The Engineering airlock’s external door was tightened to facilitate opening and closing. It still needs to be slightly lifted when opening.

· A snake (oval head, sand coloured skin) was found in the tunnel to the Science Dome at 6:00 PM and run away when it saw the crew.

Summary of GreenHab operations: The hot weather required more water for the plants.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: sand accumulated in the past EVAs was cleaned and drones were cleaned and repaired.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Burnable and Non Burnable trash is now in the Engineering airlock.